OpenStudy (anonymous):

if f(g(x))= esp(7cos(x)) what is f(x) and g(x)?

OpenStudy (amistre64):

esp? or exp?

OpenStudy (anonymous):

it says esp?

OpenStudy (amistre64):

7cos(x) is inside, so id say thats g(x)

OpenStudy (amistre64):

esp(x) would then be f(x)

OpenStudy (amistre64):

whatever esp is

OpenStudy (amistre64):

the 7 is really not critical, it can go in either place

OpenStudy (anonymous):

i mean exp sorry

OpenStudy (amistre64):

exp(x) is another way to write e^x

OpenStudy (amistre64):

f(x) = e^(7x) or e^x g(x) = cos(x) or 7 cos(x)

OpenStudy (anonymous):

it is the function for the espy awards